# FBSDiff: Plug-and-Play Frequency Band Substitution of Diffusion Features for Highly Controllable Text-Driven Image Translation

### v1

@@ III-C Implementation Details
We use the pre-trained Stable Diffusion v1.5 as the backbone diffusion model and set the classifier-free guidance scale $\omega=7.5$. We use 1000-step DDIM inversion to ensure high-quality reconstruction, i.e., $T_{inv}$=1000, and use 50-step DDIM sampling for both the reconstruction and sampling trajectory, i.e., $T$=50. Along the sampling trajectory, we allocate 55% time steps to the calibration phase and the remaining 45% steps for the non-calibration phase, i.e., $\lambda$=0.45. For the default DCT masking thresholds used in the FBS layer, we set $th_{lp}$=80 for the low-frequency band substitution (low-FBS); $th_{hp}$=5 for the high-frequency band substitution (high-FBS); $th_{mp1}$=5, $th_{mp2}$=80 for the mid-frequency band substitution (mid-FBS). The complete algorithm of FBSDiff is presented in Alg. 1.
@@ IV Experiments
Figure 4: Qualitative results of our method with different types of frequency band substitution. For low-frequency band substitution (low-FBS), the generated image is controlled by the reference image in terms of image appearance and layout; for high-frequency band substitution (high-FBS), the reference image controls the image contours of the generated image; as for mid-frequency band substitution (mid-FBS), only image layout of the generated image is controlled by the reference image. Better viewed with zoom-in.
Figure 5: Comparison among different reference image control effects achieved by low-FBS, mid-FBS, and high-FBS. Low-FBS controls image appearance and layout, high-FBS controls image contours, and mid-FBS controls only image layout.
@@ IV-A Qualitative Results
Example text-driven I2I translation results of our method are shown in Fig. 4. Our method effectively decomposes different guiding factors of the reference image to the generated image by dynamically transplanting different types of DCT frequency bands of diffusion features. The low-FBS transfers low-frequency information of the reference image into the sampling trajectory, producing the generated image that inherits the original image appearance and layout. In the mode of high-FBS that dynamically transplants high-frequency components of diffusion features, the generated image is aligned with the reference image in high-frequency contours while the low-frequency appearance is not restricted. The mid-FBS mainly imposes image layout control to the generated image since the lower-frequency appearance information and the higher-frequency contour information of the reference image are filtered out in the DCT domain. For all three types of frequency band substitution, the image translation results exhibit high visual quality and high text fidelity for both real-world and artistic-style reference images.
@@
We qualitatively compare our method with SOTA text-driven I2I translation methods including Plug-and-Play (PAP) [21], Null-text Inversion (Null-text) [16], Pix2Pix-zero [19], InstructPix2Pix (InsPix2Pix) [9], Prompt Tuning Inversion (PT-inversion) [17], StyleDiffusion [18], and VQGAN-CLIP (VQCLIP) [2], results are displayed in Fig. 7. The top panel of Fig. 7 shows that our method with low-FBS achieves better appearance consistency between the reference image and the translated result than related approaches, and is thus better suited to image creation scenario which favors inheriting the appearance and style from an existing image as much as possible. The bottom panel of Fig. 7 shows that existing SOTA text-driven I2I methods struggle at producing I2I results with large appearance change from the reference image, while our method with high-FBS excels in generating images with significantly different appearance, and is thus more suitable to image creation scenario where appearance divergence is pursued. Among the compared approaches, our method is the only one that enables flexible control over different guiding factors of the reference image, and is also the only approach that simultaneously dispenses with model training, fine-tuning, online optimization, and attention modulations.
@@
The importance of frequency band substitution (FBS) for reference image control is clearly shown in Fig. 8, from which we see that low-FBS establishes appearance and layout correlations between the reference and the generated images, while removing frequency band substitution leads to results with completely no correlation with the reference images. Moreover, as Fig. 9 displays, our method robustly adapts to varying degrees of semantic gap between the reference image and the target text prompt. The translated image of our method can still comply with the target text accurately with satisfying visual quality even in the case of very large image-text semantic discrepancy.
Besides the controllability in the guiding factors of the reference image, our method also allows continuous control over the guiding intensity simply by modulating the bandwidth of the substituted frequency band. Results displayed in Fig. 10 demonstrate the image appearance and layout guiding intensity control of our method by adjusting the low-pass filtering threshold $th_{lp}$ in the mode of low-FBS. Enlarging the value of $th_{lp}$ widens the bandwidth of the transplanted frequency band and thus increases the amount of the low-frequency guiding information of the reference image, leading to the translated image with more resemblance to the reference image. Conversely, lowering the value of $th_{lp}$ narrows the bandwidth of the substituted frequency band, which reduces the amount of the low-frequency guiding information and thus brings more variations between the reference image and the translated result.
Likewise, results in Fig. 11 demonstrate the image contour guiding intensity control of our method by adjusting the mid-pass filtering upper bound threshold $th_{mp2}$ in the mode of mid-FBS. When increasing the value of $th_{mp2}$, more high-frequency components of the reference image, namely more high-frequency guiding information, are included in the transplanted frequency band and transferred to the sampling trajectory, which results in more consistent image contours between the reference image and the translated image. On the contrary, decreasing the value of $th_{mp2}$ shrinks the transplanted high-frequency guiding information and thus leads to weaker image contour consistency.
@@ IV-B Ablation Study
To verify the rationality and effectiveness of our proposed method, we also explore and compare with other designs of frequency band substitution, including substituting the frequency band only once at $\lambda T$ time step rather than along the whole calibration phase (which we denote as Once Substitution) and substituting the full DCT spectrum rather other only a partial frequency band of it (which we refer to as Full Substitution).
The image translation results of different designs of frequency band substitution (FBS) are displayed in Fig. 12. It shows that Once Substitution produces severely noisy results rather than reasonable images, which indicates that step-by-step FBS along the whole calibration phase is of crucial importance for smooth and stable information fusion. Since image content is basically formed at the early stage of the diffusion sampling process, removing per-step feature calibration of FBS in the early sampling stage will inevitably lead to large divergence between the sampling trajectory and the reconstruction trajectory. In this case, substituting a frequency band at the intermediate time step $\lambda T$ will cause completely incoherent 2D DCT spectrum, and thus leads to abnormal image translation results after converting the diffusion features back to the spatial domain.
Besides, it also shows that Full Substitution fails to manipulate the reference image as per the text prompt. This is because substituting the full DCT spectrum is equivalent to complete feature replacement, which makes the sampling trajectory totally the same as the reconstruction trajectory during the calibration phase, the early part of the diffusion sampling process that dominates the forming of image content. Therefore, the generated image content is forced to be the same as the reference image after the calibration phase and is difficult to be modified noticeably during the subsequent non-calibration phase, the latter part of the diffusion sampling process that focuses on refining fine-grained image details rather than coarse-grained image content. Thus, the sampling results of Full Substitution closely resembles the reference image, lacking editability and text fidelity.
@@ IV-C Quantitative Evaluations
To quantitative method evaluation, we evaluate methods separately on the text-driven I2I translation task that pursues image appearance consistency and the task that pursues image appearance divergence. For the former task, we assess models’ appearance and layout preservation ability by measuring Structure Similarity ($\uparrow$), Perceptual Similarity ($\uparrow$), and Style Distance ($\downarrow$) between the reference image and the translated image pair. For the latter task, we assess models’ contour preservation and appearance alteration capabilities by measuring Structure Similarity ($\uparrow$) and Style Distance ($\uparrow$) between I2I translation pairs. For Structure Similarity measurement, we use DINO-ViT self-similarity distance [7] as the metric for Structure Distance between two images, and define Structure Similarity as 1 - Structure Distance. We use LPIPS [47] metric to measure Perceptual Similarity, and use AdaIN style loss [48] to measure Style Distance between I2I pairs. Besides, CLIP Similarity ($\uparrow$) metric is used to measure semantic consistency between the target text prompt and the translated image, i.e., text fidelity of the I2I translation results. Finally, we evaluate Aesthetic Score ($\uparrow$) of the translated images via the pre-trained LAION Aesthetics Predictor V2 model.
We sample reference images from the LAION Aesthetics 6.5+ dataset for quantitative evaluation. For the above-mentioned two tasks, we separately sample 500 reference images for each task and manually design 2 editing text prompts for each reference image, resulting in 1000 evaluation samples (pairs of reference image and target text) for each task. For evaluation of our method, we use low-FBS to handle the task pursuing appearance consistency and use high-FBS for the task pursuing appearance divergence. The average values of all the evaluation metrics are reported in Tab. I. Our method achieves top rankings for all the metrics in both two tasks, indicating superiority of our method in layout and appearance preservation with low-FBS, as well as simultaneous contour preservation and appearance modification with high-FBS. Moreover, the competitive results in CLIP Similarity and Aesthetic Score reflect that our method can generate I2I translation results with high text fidelity and visual quality.

### v2

@@ III-C Implementation Details
We use the pre-trained Stable Diffusion v1.5 as the backbone diffusion model and set the classifier-free guidance scale $\omega=7.5$. We use 1000-step DDIM inversion to ensure high-quality reconstruction, i.e., $T_{inv}$=1000, and use 50-step DDIM sampling for both the reconstruction and sampling trajectory, i.e., $T$=50. Along the sampling trajectory, we allocate 55% time steps to the calibration phase and the remaining 45% steps for the non-calibration phase, i.e., $\lambda$=0.45. For the default DCT masking thresholds used in the FBS layer, we set $th_{lp}$=80 for low-frequency band substitution (low-FBS); $th_{hp}$=5 for high-frequency band substitution (high-FBS); $th_{mp1}$=5, $th_{mp2}$=80 for mid-frequency band substitution (mid-FBS). The complete algorithm of FBSDiff is presented in Alg. 1.
@@ IV Experiments
Figure 4: Qualitative results of our method with different types of frequency band substitution. For low-frequency band substitution (low-FBS), the generated image is controlled by the reference image in terms of image appearance and layout; for high-frequency band substitution (high-FBS), the reference image controls image contours of the generated image; as for mid-frequency band substitution (mid-FBS), only image layout of the generated image is controlled by the reference image. Better viewed with zoom-in.
Figure 5: Comparison among different reference image control effects achieved by low-FBS, mid-FBS, and high-FBS. Low-FBS controls image appearance and layout, mid-FBS controls only image layout, and high-FBS controls image contours.
@@ IV-A Qualitative Results
Example text-driven I2I translation results of our method are shown in Fig. 4. Our method effectively decomposes different guiding factors of the reference image by dynamically transplanting different types of DCT frequency bands of diffusion features. The low-FBS transfers low-frequency information of the reference image into the sampling trajectory, producing translated image that inherits the original image appearance and layout. In the mode of high-FBS that dynamically transplants high-frequency components of diffusion features, the generated image is aligned with the reference image in high-frequency contours while the low-frequency appearance is not restricted. Results of mid-FBS maintain only overall image layout of the reference image, since the lower-frequency appearance information and higher-frequency contour information of the reference image are filtered out in the DCT domain. For all three modes of frequency band substitution, the image translation results exhibit high visual quality and high text fidelity for both real-world and artistic-style reference images.
@@
We qualitatively compare our method with SOTA text-driven I2I translation methods including Plug-and-Play (PAP) [21], Null-text Inversion (Null-text) [16], Pix2Pix-zero [19], InstructPix2Pix (InsPix2Pix) [9], Prompt Tuning Inversion (PT-inversion) [17], StyleDiffusion [18], and VQGAN-CLIP (VQCLIP) [2], results are displayed in Fig. 7. The top panel of Fig. 7 shows that our method with low-FBS achieves better appearance consistency between the reference image and the translated result than related approaches, and is thus better suited to image creation scenario which favors inheriting the appearance and style from an existing image. The bottom panel of Fig. 7 shows that existing SOTA text-driven I2I methods struggle at producing I2I results with large appearance change from the reference images, while our method with high-FBS excels in generating images with significantly different appearance, and is thus more suitable to image creation scenario where appearance divergence is pursued. Among the compared approaches, our method is the only one that enables flexible control over different guiding factors of the reference image, and is also the only approach that simultaneously dispenses with model training, fine-tuning, online optimization, and attention modulations.
@@
The importance of frequency band substitution (FBS) for reference image control is clearly shown in Fig. 8, from which we see that low-FBS establishes appearance and layout correlations between the reference and the generated images, while removing frequency band substitution leads to results without any correlation to the reference images. Moreover, as Fig. 9 displays, our method robustly adapts to varying degrees of semantic gap between the reference image and the target text prompt. The translated image of our method can still comply with the target text accurately with satisfying visual quality even in the case of very large image-text semantic discrepancy.
Besides the controllability in the guiding factors of the reference image, our method also allows continuous control over the guiding intensity simply by modulating the bandwidth of the substituted frequency band. Results displayed in Fig. 10 demonstrate the image appearance and layout guiding intensity control of our method by adjusting the low-pass filtering threshold $th_{lp}$ in the mode of low-FBS. Enlarging the value of $th_{lp}$ widens the bandwidth of the transplanted low-frequency band and thus increases the amount of guiding information of the reference image, leading to the translated image with more resemblance to the reference image. Conversely, lowering the value of $th_{lp}$ narrows the bandwidth of the substituted frequency band, which reduces the amount of guiding information and thus brings more variations to the translated result as compared with the reference image.
Likewise, results in Fig. 11 demonstrate the image contour guiding intensity control of our method by adjusting the mid-pass filtering upper bound threshold $th_{mp2}$ in the mode of mid-FBS. When increasing the value of $th_{mp2}$, more high-frequency components of the reference image (high-frequency guiding information) are included into the transplanted frequency band and transferred to the sampling trajectory, which results in more consistent image contours between the reference image and the translated image. On the contrary, decreasing the value of $th_{mp2}$ shrinks the transplanted high-frequency guiding information and thus leads to weaker image contour consistency.
@@ IV-B Ablation Study
To verify the rationality and effectiveness of our proposed method, we also explore and compare with other designs of frequency band substitution, including substituting the frequency band only once at $\lambda T$ time step rather than along the whole calibration phase (which we denote as Once Substitution), and substituting the full DCT spectrum rather other only a partial frequency band of it (which we refer to as Full Substitution).
The image translation results of different designs of frequency band substitution (FBS) are displayed in Fig. 12. It shows that Once Substitution produces severely noisy results rather than reasonable images, which indicates that step-by-step FBS along the whole calibration phase is of crucial importance for smooth and stable information fusion. Since image content is basically formed in the early stage of the diffusion sampling process, removing per-step feature calibration of FBS in the early sampling process will inevitably lead to large deviation of the sampling trajectory against the reconstruction trajectory. In this case, substituting a frequency band at an intermediate time step will cause completely incoherent 2D DCT spectrum, and thus leads to abnormal image translation results after converting the diffusion features back to the spatial domain.
Besides, it also shows that Full Substitution fails to manipulate the reference image as per the text prompt. This is because substituting the full DCT spectrum is equivalent to complete feature replacement, which makes the sampling trajectory totally the same as the reconstruction trajectory during the calibration phase, the early part of the diffusion sampling process that dominates the forming of image content. Therefore, the generated image content is forced to be the same as the reference image after the calibration phase and is difficult to be modified noticeably during the subsequent non-calibration phase, the latter part of the diffusion sampling process that focuses on refining fine-grained image details rather than coarse-grained image content. Thus, the sampling results of Full Substitution closely resemble the reference images, lacking editability and text fidelity.
@@ IV-C Quantitative Evaluations
For quantitative method evaluation, we separately evaluate methods on the text-driven I2I translation task pursuing image appearance consistency and the task pursuing image appearance divergence. For the former task, we assess models’ appearance and layout preservation ability by measuring Structure Similarity ($\uparrow$), Perceptual Similarity ($\uparrow$), and Style Distance ($\downarrow$) between the reference image and the translated image pair. For the latter task, we assess models’ contour preservation and appearance alteration capabilities by measuring Structure Similarity ($\uparrow$) and Style Distance ($\uparrow$) between I2I translation pairs. For Structure Similarity measurement, we use DINO-ViT self-similarity distance [7] as the metric for Structure Distance between two images, and define Structure Similarity as 1 - Structure Distance. We use LPIPS [47] metric to measure Perceptual Similarity, and use AdaIN style loss [48] to measure Style Distance between I2I pairs. Besides, CLIP Similarity ($\uparrow$) metric is used to measure semantic consistency between the target text prompt and the translated image, i.e., text fidelity of the I2I translation results. Finally, we evaluate Aesthetic Score ($\uparrow$) of the translated images via the pre-trained LAION Aesthetics Predictor V2 model.
We sample reference images from the LAION Aesthetics 6.5+ dataset for quantitative evaluation. For the above-mentioned two tasks, we separately sample 500 reference images for each task and manually design 2 editing text prompts for each reference image, resulting in 1000 evaluation samples (reference image and target text pairs) for each task. For evaluation of our method, we use low-FBS for the task pursuing appearance consistency and use high-FBS for the task pursuing appearance divergence. The average values of all the evaluation metrics are reported in Tab. I. Our method achieves top rankings for all the metrics in both two tasks, indicating superiority of our method in layout and appearance preservation with low-FBS, as well as simultaneous contour preservation and appearance modification with high-FBS. Moreover, the competitive results in CLIP Similarity and Aesthetic Score reflect that our method can generate I2I translation results with high text fidelity and visual quality.